OpenStudy (anonymous):

Give an example of a function using the variable x and explain how we evaluate a function for a given value of x

OpenStudy (mr.math):

Take a function f that represents the area of a square with a side's length x. We can write: \[f(x)=x^2.\]

OpenStudy (mr.math):

If we have a square whose side is 4 cm, we can find its area by plugging into the function: \[\text{area}=f(4)=(4\text{cm})^2=16\text{cm}^2.\]